Verify following is an identity.
(1-cos(x))/(1+cos(x))=(csc(x)-cot(x))^2

Answer by Alan3354(69443)

Verify following is an identity.
(1-cos(x))/(1+cos(x))=(csc(x)-cot(x))^2
-----------------
Change everything to sines and cosines
(1-cos(x))/(1+cos(x))=(1/sin(x) - cos(x)/sin(x))^2
(1 - cos)/(1+cos) = ((1 - cos)/sin)^2
1/(1+cos) = (1-cos)/sin^2
sin^2 = 1-cos^2
I'm convinced.